Coal is carried from a mine in West Virginia to a power plant in New York in hopper cars on a long train. The automatic hopper car loader is set to put 74 tons of coal into each car. The actual weights of coal loaded into each car are normally distributed, with mean μ = 74 tons and standard deviation σ = 0.8 ton.

(A)What is the probability that one car chosen at random will have less than 73.5 tons of coal? (Round your answer to four decimal places.)

(b) What is the probability that 42 cars chosen at random will have a mean load weight x of less than 73.5 tons of coal? (Round your answer to four decimal places.)

(c) Suppose the weight of coal in one car was less than 73.5 tons. Would that fact make you suspect that the loader had slipped out of adjustment?
Answer #1

Answer:

Given,

Mean = 74

Standard deviation = 0.8

a)

P(X < 73.5) = P((x-u)/s < (73.5 - 74)/0.8)

= P(z < -0.625)

= 0.2659855 [since from z table]

= 0.2660

b)

sample n = 42

P(X < 73.5) = P((x-u)/(sqrt(n)) < (73.5 - 74)/(0.8/sqrt(42)))

= P(z < -4.05)

= 0.0000256 [since from z table]

= 0.00003

c)

Here we can say no i.e., 0.2660 > 0.05